\begin{thebibliography}{99}
%\bibitem{ASi}
%M.Aizenman, B.Simon, Brownian motion and Harnack's inequality for 
%Schr\"odinger operators, {\em Comm. Pure Appl. Math}. {\bf 35} (1982), 209-271.

%\bibitem[Be94]{Be94} M. Beals, Optimal $L^\infty$ decay for solutions to the
%wave equation with a potential,  {\em Comm. P.D.E.} {\bf 19}(1994), no. 7-8, 1319-1369.

\bibitem{BS93}
M.~Beals, W.~Strauss, $L^p$ estimates for the wave equation 
 with a potential. 
{\em Comm. P.D.E.} {\bf 18} (1993), no. 7-8, 1365--1397.


\bibitem{BZ} J.~Benedetto, S.~Zheng, Besov spaces for the 
Schr\"odinger operator with barrier potential, to appear in {\em Complex Analysis and Operator Theory}, Birkh\"auser.
%{\em http://lanl.arXiv.org/math.CA/0411348}.

%\bibitem[BerSh]{BerSh} F.A.~Berezin, M.A.~Shubin, {\em The Schr\"odinger Equation},
%English edition, Kluwer Academic Publishers, the Netherlands 1991

\bibitem{BL76} J.~Bergh, J.~L\"ofstr\"om, 
{\em Interpolation Spaces}. Springer-Verlag, %Berlin/Heidelberg/New York, 
1976.

%\bibitem{BPST03}
 %Burq, Nicolas; Planchon, Fabrice; Stalker, John G.; Tahvildar-Zadeh, A. Shadi,
  %Strichartz estimates for the wave
 %and Schr\"odinger equations with the inverse-square potential. J. Funct. Anal. 203 (2003), no. 2, 519--549.
 
 %\bibitem{BPST04}
 %Burq, Nicolas; Planchon, Fabrice; Stalker, John G.; Tahvildar-Zadeh, A. Shadi,
  %Strichartz estimates for the wave and Schr\"odinger equations with potentials of critical decay. Indiana Univ. Math. J. 53 (2004), no. 6, 1665--1680. 
 
\bibitem{Ca63} A.~Calder\'on, Intermediate spaces and interpolation.
{\em Stud. Math.}, Seria Specjalna, %(special series)
{\bf 1} (1963), 31--34.

\bibitem{Ca64}\bysame, %A.~Calder\'on, 
Intermediate spaces and interpolation, the complex method. 
{\em Studia  Math}. {\bf 24} (1964), 113--190. 

%\bibitem{CLY} 
%S.Y.Cheng, P.Li, S.T.Yau, On the upper estimate of the heat kernel of
%a complete Riemannian manifold, {\em Amer. J. Math.} {\bf 103} (1981), 1021-1063.


%\bibitem{CS88} M. Christ, C.D. Sogge, The weak type $L^1$ convergence of eigenfunction expansions for pseudodifferential operators, Invent. Math. 94 (2) (1988) 421--453.

%\bibitem{CMS86} M.Cwikel, M.Milman,Y.Sagher, Complex interpolation of some quasi-Banach spaces, {\em J. Funct. Anal}. {\bf 65} (1986), 339-347

%\bibitem{CouS06} T.~Coulhon, A.~Sikora, 
%Gaussian heat kernel upper bounds via Phragm\'en-Lindel\"of theorem. http://xxx.lanl.gov/abs/math/0609429.

%\bibitem{CZh05}
%T.Coulhon, Q. Zhang, Heat kernels on forms.  arXiv 2005.

%\bibitem{CGM02} M.~Cowling, A.~Sikora, A spectral multiplier theorem for a sublaplacian
%on $SU(2)$,  Math. Z. 238 (2001), no. 1, 1--36.  

%\bibitem{Cu00}
%S.~Cuccagna, On the wave equation with a potential, {\em Comm. P. D. E}. {\bf 25} (2000), 1549--1565.

\bibitem{CuS01}
S.~Cuccagna, P.~Schirmer, On the wave equation with a magnetic potential. 
{\em Comm. Pure Appl. Math}. {\bf 54} (2001) no. 2, 135--152.

%\bibitem{DF05} P.~D'Ancona, L.~Fanelli, Decay estimates for the wave and 
%Dirac equations with a magnetic potential. http://arXiv.org/abs/math/0504370.

\bibitem{DP05} P.~D'Ancona, V.~Pierfelice, 
On the wave equation with a large rough potential.  {\em J. Funct. Anal}. 
{\bf 227} (2005), no. 1, 30--77.
%http://arXiv.org/math.AP/0310199.

\bibitem{Da89}
E.~Davies, {\em Heat Kernels and Spectral Theory}. Cambridge University Press, Cambridge, 1989. %MR0990239 (90e:35123) 

%\bibitem{Da89b} \bysame,
%Pointwise bounds on the space and time derivatives of heat kernels. {\em J. Operator Theory}
%{\bf 21} (1989), 367--378.

\bibitem{Da95}\bysame, Uniformly elliptic operators with measurable coefficients. {\em J. Funct. Anal}. {\bf 132} (1995), no.1, 141--169.

%\bibitem{DP89} Davies, E. B.; Pang, M. M. H.,  Sharp heat kernel bounds for some Laplace operators.  
%Quart. J. Math. Oxford Ser. (2) 40 (1989), no. 159, 281--290.

%\bibitem{DS91}
%Davies, Simon,  Heat kernel for noncritical Schrodinger semigroup.  JFA 1991.

%\bibitem{MM87}
%L. De-Michele, G. Mauceri, $H^p$ multipliers on stratified groups, Ann. Mat. Pura Appl. (4) 148 (1987) 353--366. 

%\bibitem{DL79}
%Harold Donnelly and Peter Li, Pure point spectrum and negative curvature for noncompact manifolds,
% Duke Math. J.  {\bf 46} (1979), no.3, 497Ð503.

%\bibitem{DSch}
%N.~Dunford, J.~Schwartz, {\em Linear Operators}. Vol. I, Interscience Publishers, London, 1958.

%\bibitem{DM99} X.T.~Duong, A.~McIntosh, 
%Singular integral operators with non-smooth kernels on irregular domains,
%Rev. Mat. Iberoamericana,  {\bf 15} (1999), no.2: 233-265.

\bibitem{DOS02} X.~Duong, E.~Ouhabaz, A.~Sikora, Plancherel type estimates and sharp spectral multipliers.
 {\em J. Funct. Anal}. {\bf 196} (2002),  443--485.

%\bibitem{DY05}
% X.-T.~Duong; L.~Yan, Duality of Hardy and BMO spaces associated with operators with heat kernel bounds.
 %{\em J. Amer. Math. Soc}. {\bf 18} (2005), 943-973. 

\bibitem{D97} J.~Dziuba\'nsk, Triebel-Lizorkin spaces associated with 
Laguerre and Hermite expansions. {\em Proc. Amer. Math. Soc.} {\bf
125} (1997), 3547--3554.


%\bibitem{D99}\bysame,  Spectral multiplier theorem for $H^1$ spaces
%associated with some Schr\"odinger operators, {\em Proc. A.M.S.}  {\bf 127} (1999), no. 12, 3605-3613.

%\bibitem{DZ05} J.~Dziuba\'nski, J.~Zienkiewicz, 
%Hardy spaces $H\sp 1$ for Schr\"odinger operators with compactly supported potentials. 
%{\em Ann. Mat. Pura Appl}. (4) 184 (2005), no. 3, 315--326.

%\bibitem{DGMTZ05}
%J. Dziubanski, GUSTAVO GARRIG\'oS, T. Martínez, J.L. Torrea, J. Zienkiewicz, 
%"BMO spaces related to Schrodinger operators with potentials satisfying a reverse Holder inequality", 
%Math. Z. 249, 329-356 (2005).

%\bibitem[DZ99]{DZ99} J.~Dziuba\'nski and J.~Zienkiewicz, Hardy space $H^1$
%associated to Schr\"odinger operator with potential satisfying
%reverse H\"older inequality, {\em Rev. Mat. Iberoamericana} {\bf
%15}(2)( 1999), 279-296.

%\bibitem{DZ97} %littlewood-paley
%Hardy spaces associated with some Schrodinger operators, J. Dziubanski and J. Zienkiewicz
%Studia Math. 126 (1997), 149-160.

%\bibitem{E95} J.~Epperson, Triebel-Lizorkin Spaces for Hermite
%expansions. {\em Studia Math.} {\bf 114} (1995), no.1, 87--103.

%\bibitem[E96]{E96} ---, Hermite multipliers and pseudo-multipliers, {\em
%Proc. Amer. Math. Soc.} {\bf 124} (1996), no. 7, 2061--2068.

%\bibitem{E97}---, Hermite and Laguerre wave packet expansions. {\em Studia
%Math.} {\bf 126} (1997), no. 3, 199--217.


%\bibitem{Fe70} C.~Fefferman, Inequalities for strongly singular convolution operators. {\em Acta Math}. {\bf 124} (1970), 9-36.

%\bibitem{FeS71}
%C.~Fefferman and E.~Stein, Some maximal inequalities. {\em Amer. J. Math}. {\bf 93} (1971), 107--115.

%\bibitem[FM84]{FM84}
%V.~Ferrari and B.~Mashoon, New approach to the quasinormal modes of
%a black hole. {\em Phys. Rev. D}. {\bf 30} (1984), 295-304.

%\bibitem[Flu74]{Flu74} S.~Fl\"ugge, {\em Practical Quantum Mechanics}, Springer-Verlag, 1974.

%\bibitem{FJW} M.~Frazier, B.~Jawerth, G.~Weiss. {\em Littlewood-Paley Theory
%and the Study of Function Spaces}.  Conference Board of the Math. Sci. {\bf 79}, 1991.

%\bibitem{FMV05} G. Furioli, C. Melzi, A. Veneruso, 
% Littlewood-Paley decompositions and Besov spaces on Lie groups of polynomial growth,  math.CA/0502406 (2005).

%\bibitem{GV03} 
%V.~Georgiev, N.~Visciglia, Decay estimates for the wave equation
%with potential, Comm. P.D.E. {\bf 28} (2003), %nos.7,8 
%1325-1369.

%\bibitem{G04a} M.~Goldberg, Dispersive bounds for the three-dimensional
%Schr\"odinger equation with almost critical potentials. arXiv. 2004.

\bibitem{GV95} 
J.~Ginibre, G.~Velo, Generalized Strichartz inequalities for
the wave equation. {\em J. Funct. Anal}. {\bf 133} (1995), 50-68.

%\bibitem{Gr95}
%A. GrigorÕyan, Upper bounds of derivatives of the heat kernel on an arbitrary complete manifold.
%{\em  J. Funct. Anal}. 1995.

%\bibitem{Gr97} \bysame, GAUSSIAN UPPER BOUNDS FOR THE HEAT KERNEL ON ARBITRARY MANIFOLDS, J. Differential Geometry 45 (1997) 33--52.

\bibitem{Gr66}  P.~Grisvard, Commutativit\'e de deux foncteurs d'interpolation et applications.
 {\em J. Math. Pures Appl}. {\bf 45} (1966), 143--290.

%\bibitem[GH98]{GH98} C.-A.~Guerin, M.~Holschneider, Time-dependent scattering
%on fractal measures,  {\em J. Math. Physics} {\bf 39}(8), 1998.

\bibitem{He90a} 
W.~Hebisch: A multiplier theorem for Schr\"odinger operators.
\textit{Colloq. Math}, {\bf 60/61} (1990),  no. 2, 659--664.

%\bibitem{He90b}\bysame, Almost everywhere summability of eigenfunction expansions associated to elliptic 
%operators. \textit{Studia Math} {\bf 96} (1990), no. 3, 263--275.

%\bibitem{He95}\bysame, Functional calculus for slowly decaying
%kernels, preprint, 1995.\\ 
%http://www.math.uni.wroc.pl/\symbol{126}hebisch

%\bibitem{HMM06}
%Waldemar Hebisch, Giancarlo Mauceri and Stefano Meda, 
%Spectral multipliers for sublaplacians with drift on Lie groups. preprint.

%\bibitem{He92}
%Sigurdur Helgason, Huygens' Principle for Wave Equations on Symmetric Spaces. 
%J. Funct. Anal. 107 (1992) 279-288. 

%\bibitem{Hor68} L.~H\"ormander, The spectral function of an elliptic operator, Acta Math. {\bf 121} (1968), 193-218.

%\bibitem{Hor60}\bysame, Estimates for translation invariant operators in
%$L^p$ spaces, Acta Math. {\bf 104} (1960), 93-140.

%\bibitem{Hor83}\bysame, The analysis of linear partial differential operators,
%II, Springer-Verlag, Berlin, 1983. Differential operators with constant coefficients.

%\bibitem{JK79}
%A.Jensen, T.Kato, Spectral properties of Schr\"odinger operators
%and time-decay of the wave functions, Duke Math. J. {\bf 46}
%(1979), 583-611.

%\bibitem{J84}
%A.~Jensen, Spectral properties of Schr\"odinger operators
%and time-decay of the wave functions, Results in $L_2(R^4)$, 
 %J.~Math.~Anal.~Appl. {\bf 101} (1984), 397-422.

%\bibitem{J80}
%A.~Jensen, Spectral properties of Schr\"odinger operators
%and time-decay of the wave functions, Results in $L_2(R^m)$, 
%$m\ge 5$, Duke Math. J. {\bf 47} (1980), 57-80.

\bibitem{JN94} A.~Jensen, S.~Nakamura, Mapping properties of functions
of Schr\"odinger operators between $L^p$ spaces and Besov spaces. %in {\em Spectral and Scattering Theory and Applications}, 
\emph{Advanced Studies in Pure Math}. {\bf 23} (1994), 187--209.

%\bibitem{JN01}
%A.~Jensen, G.~Nenciu, A unified approach to resolvent expansions
%at thresholds, {\em Rev. Math. Phys}. {\bf 13} (2001) no.6, 717-754.



%\bibitem[KMS]{KMS} S.~Klainerman, M.~Machedon and J.~Stalker, Decay of
%solutions to the wave equation on a spherically symmetric static  background, {\em preprint}.

%\bibitem{KR} H.Koch, F.Ricci, Spectral projections for the twisted Laplacian, arXiv, 2004. 

%\bibitem[Ko98]{Ko} L.A.~Kolasa, Oscillatory integrals with nonhomogeneous
%phase functions related to Sch\"odinger equations, {\em
%Canad. Math. Bull.}, {\bf 41}(3), 1998, 306-317.

%\bibitem{Ku00} Kurata, K.: An estimate on the heat kernel of magnetic Schr\"odinger operators and uniformly elliptic operators with non-negative potentials. J. London Math. Soc. (2) 62, 885--903 (2000) 

%\bibitem{L86}
%P.~Li,  Large time behavior of the heat equation on complete manifolds 
%with nonnegative Ricci curvature, {\em Ann. Math}. {\bf 124} (1986), 1-21.

%\bibitem{LY86}
%P.Li, S.T.Yau, On the parabolic kernel of the Schr\"odinger
%operator, Acta Math. {\bf 156} (1986), 153-201. 

%\bibitem[Le75]{Le75}
%N.N.~Lebedev, Special functions and their applications,
%Translation by R.A.~Silverman, New York, Dover Publications, 1972. 

\bibitem{LP64}
J.~Lions, J.~Peetre, Sur une classe d'espaces d'interpolation.
{\em Inst. Hautes \'Etudes Sci. Publ. Math}. {\bf 19} (1964), 5--68.

%\bibitem{Ma82} C.~Markett, Mean Ces\`aro summability of Laguerre expansions
%and norm estimates with shifted parameter, {\em Anal. Math}. {\bf 8} (1982), 19-37.

%\bibitem{MM90} G. Mauceri, S. Meda, Vector-valued multipliers on stratified groups, Rev. Mat. Iberoamericana 6 
%(3-4) (1990) 141--154.

%\bibitem{Mi65}
%S.G. Mikhlin, Multidimensional Singular Integrals and Integral Equations, Pergamon Press, Oxford, 1965 (translated from the Russian by W. J. A. Whyte. Translation edited by I. N. Sneddon). 

%\bibitem{Mu70}B.~Muckenhoupt, Mean convergence of Hermite and Laguerre series
%I, II.  {\em    Trans. A.M.S}. {\bf 147} (1970), 419-431; {\em ibid.},  {\bf 147} (1970), 433-460.

%\bibitem{MS94} D. M\"uller, E.M. Stein, On spectral multipliers for Heisenberg and 
%related groups, J. Math. Pures Appl. (9) 73 (4) (1994) 413--440. 

%\bibitem{MT}
%D. M\"uller, Ch. Thiele,  Wave equation and multiplier estimates on $ax+b$ groups  ; to appear Studia Math. 

%\bibitem{Mu82}
%M.~Murata, Asymptotic expansions in time for solutions
%of Sch\"odinger-type equations, {\em J. Funct. Anal}. {\bf 49} 
%(1982), 10-56.

%\bibitem{Nar} E.K. Narayanan, Multipliers for the twisted Laplacian, Colloq. Math.  

%\bibitem[Ob90]{Ob} F.~Oberhettinger, {\em Tables of Fourier Transforms and
%Fourier transforms of Distributions}, Springer-Verlag, 1990.    

%\bibitem[Ob73]{Ob2}---, {\em Fourier Expansions :A Collection of Formulas},
%Academic Press, New York-London, 1973.

%\bibitem[O87]{O87}
%G.~\'Olafsson, Fourier and Poisson transformation associated to a semisimple symmetric space, {\em Invent. Math}. {\bf 90} (1987), no. 3, 605--629.

\bibitem{OZ06}
G.~\'Olafsson, S.~Zheng, Function spaces associated with 
Schr\"odinger operators: the P\"oschl-Teller potential. {\em J. Fourier Anal. Appl}.  {\bf 12} (2006), no.6, 653--674. 

\bibitem{OZ08} G.~\'Olafsson, S.~Zheng,
Harmonic analysis related to Schr\"odinger operators. {\em Contemporary Mathematics}
{\bf 464} (2008), 213--230.

%\bibitem{OS92}
%G.~\'Olafsson, H. Schlichtkrull, Wave propagation on Riemannian symmetric 
%space. J. Funct. Anal. 107 (1992) 270-278.

\bibitem{ou06}
E.~Ouhabaz, Sharp Gaussian bounds and $L^p$-growth of semigroups associated with
elliptic and Schr\"odinger operators.  {\em Proc. A.M.S}. {\bf 134} (2006), no. 12, 3567--3575. 

\bibitem{Pe76} J.~Peetre, {\em New Thoughts on Besov Spaces}. Duke Univ. Math. Series I. Durham: Duke Univ. 1976.

\bibitem{Pe63}\bysame, %J.~Peetre, 
A theory of interpolation of normed spaces. Notes, Brasilia, 1963 
(Notas de Matematica, no.39, 1968).


%\bibitem{Pe75} J.~Peetre,  On space of Triebel-Lizorkin type, Ark. Math. 13 (1975), 123-130. 


%\bibitem{Sa79} K.~Saka, Besov spaces and Sobolev spaces on a nilpotent Lie group, {\em T$\hat{o}$hoku Math. J.} {\bf 31} (1979), 383-437. 

%\bibitem[Sch]{Sch} M.~Schechter, {\em Spectra of Partial Differential Operators},
%North-Holland series in applied math. and mechanics, {\bf 14}, 1986.

%\bibitem{Sch05b} 
%\bysame, A remark on Littlewood-Paley theory for the distorted  Fourier transform, {\em http://lanl.arXiv.org/math.AP/0508577}, (2005) 

%\bibitem[Sch05c]{Sch05c} 
%\bysame, Spectral theory and nonlinear PDE: A survey. 
%{\em http://lanl.arXiv.org/math.AP/0509019}, (2005).

\bibitem{SchmTr87} H.-J. Schmeisser, H.~Triebel, {\em Topics in Fourier
Analysis  and Function spaces}. Wiley-Interscience, 1987.

%\bibitem{ScTr86}  H.-J. Schmeisser, H. Triebel, Topics in Fourier Analysis and Function Spaces, John Wiley \& Sons (1987).


\bibitem{Si82}
B.~Simon, Schr\"odinger semigroups. {\em Bull. Amer. Math. Soc}. {\bf 7}
(1982), no.3, 447--526.

%\bibitem{Sk03}
%L.~Skrzypczak, Heat extensions, optimal atomic decompositions and Sobolev embeddings in presence of symmetries on manifolds.  
%Math. Z. 243 (2003), no. 4, 745--773.

%\bibitem{Sk98b} L.~Skrzypczak, %\bysame, 
%The Triebel-Lizorkin scale of function spaces for the Fourier-Helgason transform.  {\em Math. Nachr}.  {\bf 190}  (1998), 251-274. 

%\bibitem{Sk99b} \bysame,
%Vector-valued Fourier multipliers on symmetric spaces of the noncompact type. 
%Monatsh. Math. 119 (1995), no. 1-2, 99--123.

%\bibitem{So87} 
%C.D. Sogge, On the convergence of Riesz means on compact manifolds, Ann. of Math. (2) 126 (2) (1987) 439--447. 


%\bibitem{Str01}
%R.S. Strichartz, Function spaces on fractals, preprint, 2001. 

%\bibitem{St70} E.~Stein, {\em Topics in Harmonic Analysis Related to the Littlewood-Paley Theory}. Princeton University Press, Princeton NJ, 1970 

%\bibitem{St93} E.~Stein, {\em Harmonic analysis, Real-variable methods, orthogonality, and oscillatory integrals}, Princeton Univ. Press, 1993

\bibitem{Taib} M.~Taibleson, On the theory of Lipschitz spaces of distributions on Euclidean $n$-spaces, I,II.
{\em J. Math. Mech}. {\bf 13} (1964), 407--479; {\bf 14}   (1965), 821--839.

%\bibitem[Tay96]{Tay2} M.~Taylor, Partial Differential Equations II, {\em Qualitative
%Studies of Linear Equations}, Applied Mathematical Sciences {\bf 116},
%Springer Verlag, 1996.

%\bibitem[Ti58]{Ti58} E.C.~Titchmarsh, {\em Eigenfunction Expansions Associated with
%Second-Order Differential Equations. (Part I, II) }, Oxford Univ. Press, 1958.


\bibitem{Tr78} H.~Triebel,
{\em Interpolation Theory, Function Spaces, Differential Operators}. 
%Berlin, VEB Deutsch. Verl. Wissenschaften 1978, 
Amsterdam, North-Holland,  1978.

\bibitem{Tr83} H.~Triebel, {\em Theory of Function Spaces}.
Birkh\"{a}user, 1983.

%\bibitem{Tr92} ---, {\em Theory of Function Spaces II}, Monographs Math. {\bf 84}, Birkh\"auser, Basel, 1992.

%\bibitem{Voi86}
%J.~Voigt, Absorption semigroups, their generators, and Schrodinger semigroups,
%{\em J. Funct. Anal}.  {\bf 67} (1995), 551-581.


\bibitem{Y95}
K.~Yajima, The $W^{k,p}$-continuity of wave operators for Schr\"odinger 
operators. {\em J. Math. Soc. Japan} {\bf 47} (1995), 551--581.

%\bibitem{Y95b}  Yajima, K., The $W^{k,p}$-continuity of wave operators for Schr\"odinger operators III. J. Math. Sci. Univ. Tokyo 2, 311--346 (1995)

%\bibitem{Y99} \bysame, $L^p$-boundedness of wave operators for two dimensional Schr\"odinger operators,  Comm.Math. Phys. {\bf 208} (1999), 125-152. 


%\bibitem[Zh01]{Zha} Q.~Zhang, Global bounds of Schr\"odinger heat kernels
%with negative potentials, {\em J. Func. Anal.} {\bf 182} (2001), no.2, 344-370. 

% \bibitem{Zh00} Zhang, Qi S. Large time behavior of Schr\"odinger heat kernels and applications. Comm. Math. Phys. 210 (2000), no. 2, 371--398. (Reviewer: Thierry Coulhon)
 

%\bibitem[Z06a]{Z06a}---, Perturbed Fourier transform  associated to Schr\"odinger operators, {\em preprint} (2006).

\bibitem{Z06a} 
S.~Zheng,  Littlewood-Paley theorem for  Schr\"odinger operators. 
{\em Anal. Theory. Appl}. {\bf 22} (2006), no.4, 353--361.

%\bibitem{Z06b}\bysame, Spectral multipliers for Schr\"odinger operators I.  Preprint. 

%\bibitem{Z08}\bysame, Spectral calculus, function spaces and dispersive equations with a critical potential. Preprint. 

%\bibitem{Z07}\bysame, {Spectral calculus, function spaces and dispersive equation with a critical potential}. In preparation.

%\bibitem[Z05]{Z05} ---, {\em Littlewood-Paley theory, atomic decomposition and Schr\"odinger equations on $\R^d$},
% Presented at the 111$^{\text{th}}$ AMS Conference, Atlanta, Jan. 2005.
\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{amsmath,amssymb,amsthm,amscd,amsopn,amsfonts,amsxtra} 
\usepackage{latexsym,array,verbatim,url}
\usepackage[mathscr]{eucal}
\usepackage{graphicx,psfrag,epsfig,color}

\usepackage[small,nohug,heads=littlevee]{diagrams} 
\diagramstyle[labelstyle=\scriptstyle]

\newtheorem{theorem}[subsection]{Theorem} 
\newtheorem{lemma}[subsection]{Lemma}

\newtheorem{assumption}[subsection]{Assumption}

\newtheorem{definition}[subsection]{Definition}

\def\ga{\gamma}

\def\veps{\varepsilon}
\def\vphi{\varphi}
\def\De{\Delta}
\def\Ga{\Gamma}

\def\supp{\mathrm{supp}}

\def\cL{\mathcal{L}}
\def\inv{^{-1}}
\def\iy{\infty}
\def\pa{\partial} 

\newcommand{\cal}{\mathcal}

\newcommand{\Z}{\mathbb{Z}}
\newcommand{\R}{\mathbb{R}}
\newcommand{\C}{\mathbb{C}}
\newcommand{\N}{\mathbb{N}}

\def\sideremark#1{\ifvmode\leavevmode\fi\vadjust{\vbox to0pt{\vss
 \hbox to 0pt{\hskip\hsize\hskip1em
\vbox{\hsize2cm\tiny\raggedright\pretolerance10000
 \noindent #1\hfill}\hss}\vbox to8pt{\vfil}\vss}}}%

\begin{document}
\title[Interpolation Theorems for Self-adjoint  operators]
{Interpolation Theorems for Self-adjoint  operators}
\author{Shijun Zheng}
\address{Department of Mathematical Sciences\\
Georgia Southern University\\
Statesboro, GA 30460-8093}
\email{szheng@georgiasouthern.edu}
\urladdr{http://math.georgiasouthern.edu/\symbol{126}{szheng}}
\keywords{interpolation, functional calculus}
\subjclass[2000]{Primary:  42B35, 
42B25; Secondary: 
35J10} 
\date{\today}

\begin{abstract}  
We prove  a complex and a real interpolation \mbox{theorems} on Besov spaces and Triebel-Lizorkin spaces 
associated with a selfadjoint operator $\cL$, without assuming the gradient estimate for its spectral kernel.
The result applies to the cases where $\cL$ is a uniformly elliptic operator or a Schr\"odinger operator
with electro-magnetic potential. 
\end{abstract}

\maketitle 

\section{Introduction and main result}\label{S1} 
Interpolation of 
function spaces has  
played an important role in classical Fourier analysis and PDEs \cite{BS93,CuS01,
GV95,JN94,OZ08}. 
Let $\cL$ be a selfadjoint  operator in $L^2(\R^n)$. Then, for a Borel measurable function $\phi$: $\R\to \C$, 
we define $\phi(\cL)$ using functional calculus. 
 In \cite{JN94,D97,BZ,OZ06}  several authors introduced and studied the 
Besov spaces and Triebel-Lizorkin spaces associated with Schr\"odinger operators. In this note we 
present an 
interpolation 
result on these spaces for 
$\cL$.  

Let $\{\varphi_j\}_{j=0}^\iy\subset C_0^\infty ({\R}) $ be a dyadic system satisfying 
(i)  $\supp \,\vphi_0 \subset \{ x: |x|\le 1\}$,
 $\supp\; \varphi_j
\subset \{ x: 2^{j-2}\le |x|\le 2^{j}\} $, $j\ge 1$, 
(ii) $|\vphi_j^{(k)}(x)|\le c_k 2^{-kj}$ for all $j$, $k\in \N_0=\N\cup\{0\}$,
(iii) $\displaystyle{
\sum_{j=0}^\infty |\varphi_j(x)| \approx 1, \quad \forall x.}$
Let $\alpha\in \R$,  $1\le p,q\le \infty$. 
The inhomogeneous \emph{Besov space} \emph{associated with} $\cL$, denoted by ${B}_p^{\alpha,q}(\cL)$, %
is defined to be the completion of $\mathcal{S}(\R^n)$, the Schwartz class,  with respect to  
the norm 
$$ \Vert f\Vert_{{B}_p^{\alpha,q}(\cL) } = 
\big(\sum_{j=0}^{\infty}
 2^{j\alpha q} \Vert \vphi_j(\cL)f \Vert_{L^p}^q \big)^{1/q}\, .
$$
Similarly, the inhomogeneous {\em Triebel-Lizorkin space associated with $\cL$}, denoted by ${F}_p^{\alpha,q}(\cL)$, 
is defined by the norm
\begin{equation*}\label{eq:F-norm}
\Vert f\Vert_{{F}_p^{\alpha,q}(\cL)} 
= 
\Vert \big(\sum_{j=0}^{\infty} 2^{j\alpha q} \vert \vphi_j(\cL)f \vert^q\big)^{1/q}\Vert_{L^p} \, .
\end{equation*} 

The following assumption on the kernel of $\phi_j(\cL)$ is fundamental in the study of 
function space theory. 
 Let $\phi(\cL)(x,y)$ denote the integral 
kernel of $\phi(\cL)$. 
\begin{assumption}\label{a:phi-dec-Om}  
Let  $\phi_j\in C_0^\infty(\R)$ satisfy conditions (i), (ii) above.  
Assume that 
  there exist some $\veps>0$ 
and a constant 
 $c_n>0$ such that for all $j$
\begin{equation}\label{e:ker-phi-dec}
\vert  \phi_j(\cL)(x,y)\vert 
\le c_n\frac{2^{nj/2}}{(1+2^{j/2}|x-y|)^{n+\veps}}\,. 
\end{equation}
\end{assumption}
This is the same condition assumed in \cite{Z06a,OZ08} except that we drop the 
gradient estimate condition on the kernel. 
This is the case when $\cL$ is a Schr\"odinger operator $-\De+V$, 
$V\ge 0$ belonging to $L^1_{loc}(\R^n)$ \cite{He90a,ou06} or $\cL$ is a uniformly elliptic operator in $L^2(\R^n)$
\cite[Theorem 3.4.10]{Da89}. 

In what follows, $[A,B]_\theta$ denotes the usual complex interpolation between two Banach spaces; 
$(A,B)_{\theta,r}$ the real interpolation, see \mbox{Section \ref{S2}.} The notion 
 $T$: $X\to Y$ means that the \mbox{linear} operator $T$ is bounded from $X$ to $Y$.

\begin{theorem}[complex interpolation]\label{th:F-B-com-int}
Suppose that $\cL$ is a selfadjoint operator satisfying Assumption \ref{a:phi-dec-Om}. 
Let $0<\theta<1$,  $s=(1-\theta)s_0+\theta s_1 $, $s_0,s_1\in\R$ 
and  \begin{equation*}
\frac{1}{p}=\frac{1-\theta}{p_0}+ \frac{\theta}{ p_1},\quad 
\frac{1}{q}=\frac{1-\theta}{q_0}+ \frac{\theta}{ q_1}.  \end{equation*}
(a) If  $ 1< p_i<\infty$,  $1< q_i< \infty$, $i=0, 1$, then 
$$[{F}_{p_0}^{s_0,q_0}(\cL),{F}_{p_1}^{s_1,q_1}(\cL)]_\theta={F}_{p}^{s,q}(\cL). 
$$
(b) If $ 1\le p_i\le\iy$, $1\le q_i\le\infty$, $i=0,1$, 
 then  $$[{B}_{p_0}^{s_0,q_0}(\cL),{B}_{p_1}^{s_1,q_1}(\cL)]_\theta
={B}_{p}^{s,q}(\cL). 
$$
c) If $T: {F}_{p_0}^{s_0,q_0}(\cL)\to {F}_{\bar{p}_0}^{\bar{s}_0,\bar{q}_0}(\cL)$ 
and $T: {F}_{p_1}^{s_1,q_1}(\cL)\to {F}_{\bar{p}_1}^{\bar{s}_1,\bar{q}_1}(\cL)$, 
then $T: {F}_{p}^{s,q}(\cL) \to F_{\bar{p}}^{\bar{s},\bar{q}}(\cL) $,
where $\bar{s},\bar{p},\bar{q}$ and $\bar{s}_i,\bar{p}_i,\bar{q}_i$,  
satisfy the same relations as those for $s,p,q$ and $s_i,p_i,q_i$, $1<p_i,q_i<\iy$.
 Similar statement holds for $B^{s,q}_p(\cL)$.
\end{theorem}

 Complex  interpolation method
 originally was due to Calder\'on \cite{Ca63} and Lions and Peetre \cite{LP64}; see also \cite{Gr66,Taib}. 
The classical interpolation \mbox{theory} for Besov and Triebel-Lizorkin spaces on $\R^n$ has been given systematic treatments 
in \cite{Pe76}, \cite{BL76}, and \cite{Tr78,Tr83}. 
 There are  interesting discussions on interpolation theory  
in \cite{Pe76} and \cite{Tr83, Tr78, SchmTr87} for generalized Besov spaces associated with differential operators, which 
requires certain Riesz summability for 
$\cL$ that seems a nontrivial condition to verify. Nevertheless, we would like to mention that the Riesz summability, the spectral multiplier
theorem and the decay estimate in (\ref{e:ker-phi-dec}) are actually intimately related \cite{DOS02,OZ08}.

The real interpolation result for $B^{\alpha,q}_p(\R^n)$, $F^{\alpha,q}_p(\R^n)$ can be found in 
\cite{Pe76} and \cite{Tr83,Tr78}. Following the proof as in the classical case, 
but applying the estimate in 
(\ref{e:ker-phi-dec}) in stead of  spectral multiplier result, we obtain   
\begin{theorem}[real interpolation]\label{th:F-B-real-int}
Suppose that $\cL$ 
 satisfies Assumption \ref{a:phi-dec-Om}. Let 
\mbox{$0<\theta<1$,} $1\le r\le \infty$, 
  $s=(1-\theta)s_0+\theta s_1 $,  $s_0\neq s_1$. \\
(a) If $1\le p<\iy, 1\le q_1,q_2 \le\infty$, then
$$(F_{p}^{s_0,q_0}(\cL),F_{p}^{s_1,q_1}(\cL) )_{\theta,r}=B_{p}^{s,r}(\cL). 
$$
(b) If $ 1\le p,  q_1,q_2\le\infty$, 
then  $$(B_{p}^{s_0,q_0}(\cL),B_{p}^{s_1,q_1}(\cL) )_{\theta,r}
=B_{p}^{s,r}(\cL). 
$$ 
\end{theorem}







 The homogeneous spaces $\dot{B}_p^{\alpha,q}(\cL)$ and $\dot{F}_p^{\alpha,q}(\cL)$ can be defined using $\{\varphi_j\}_{j=-\iy}^\iy$ in (i) to (iii), 
instead of  $\{\vphi_j\}_{j=0}^\iy$. Then the analogous results of Theorem \ref{th:F-B-com-int} and Theorem \ref{th:F-B-real-int} hold. 

\section{interpolation for $\cL$}\label{S2} 


Theorem \ref{th:F-B-com-int} and Theorem \ref{th:F-B-real-int} are part of 
the abstract interpolation theory for $\cL$. 
In this section we present  the outline of their proofs. 
It was mentioned   in \cite{SchmTr87} that the interpolation associated with $\cL$  
is a ``subtle and difficult" subject, which normally relies on the very property of $\cL$.

\subsection{Complex interpolation} 
The proof of Theorem \ref{th:F-B-com-int} is similar to that
 given in \cite{Tr78} in the Fourier case. 
The insight is that the three line theorem (involved in Riesz-Thorin or Calder\'on's constructive proof
for $L^p$ spaces) 
 reflects the fact that the value of an 
 analytic function 
 in the interior of a domain is 
determined by its boundary values. 
\begin{definition}
Let $(A_0,A_1)$ be an interpolation couple , i.e.,
$A_0,A_1$ are (complex) Banach spaces, linearly and continuously 
embedded in a Hausdorff space $\cal{H}$. 
The space $A_0\cap A_1$ is endowed with the norm 
$\Vert a\Vert_{A_0\cap A_1}= 
\max\{\Vert a\Vert_{A_j},j=0,1\}$. The space $A:=A_0+A_1$ 
is endowed with the norm 
\begin{equation*}
\Vert a\Vert_{A}=\inf
\{\Vert a_0\Vert_{A_0}+\Vert a_1\Vert_{A_1}: a_0\in A_0,a_1\in A_1\} .
\end{equation*}


Let $S=\{z\in\C:0\le\Re z\le 1\}$ and $\bar{S}$ its closure. 
Denote $F$ the class of all $A$
-valued functions $f(z)$ on 
$\bar{S}$ such that $z\mapsto f(z)\in A$ is analytic in $S$ and continuous on $\bar{S}$,
satisfying
\begin{enumerate}
\item[(i)]  \[
\sup_{z\in \bar{S}} \Vert f(z)\Vert_A \;\text{is finite}. \]
\item[(ii)] The mapping $t\mapsto f(j+it)\in A_j$
 are continuous from $\R$ to $A_j$, $j=0,1$.\end{enumerate}
Then $F$ is a Banach space with the norm
\[
\Vert f\Vert_F= \max_j\{\sup_t  \Vert f(j+it)\Vert_{ A_j}\}.
\]
For $0<\theta<1$ we define the interpolation space $[A_0,A_1]_\theta$
as
\[ [A_0,A_1]_\theta:=\{a\in A: \exists f\in F
\;\text{with}\; f(\theta)=a\}  .\]
Then $[A_0,A_1]_\theta$ is a Banach space equipped with the norm
\[ \Vert a\Vert_\theta :=\inf\{ \Vert f\Vert_F: f\in F\, and\, f(\theta)=a\}.\]
\end{definition}








\subsection{Outline of the proof of Theorem \ref{th:F-B-com-int}} 
Let $\{\phi_j\},\{\psi_j\}$ satisfy the conditions in (i)-(iii) and
 $\sum_j \psi_j(x)\phi_j(x)=1$.
Define the operators $S: 
f\mapsto \{\phi_j(\cL)f\}$,  and $R:g\mapsto \sum_j\psi_j(\cL)g$.
The proof for part (a) follows from  the commutative diagram
$$\begin{CD}
F^{s,q}_p(\cL)  @>S >>L^{p}(\ell^q)\\ 
@VIdVV             @VVIdV\\    
F^{s,q}_p(\cL)  @<R<<L^p(\ell^q) 
\end{CD}
$$ 
and Lemma \ref{lem:vector-Lp-int} and
Lemma \ref{lem:vector-lq-int},  
which are  interpolation results for Banach space valued $L^p$ and $\ell^q$ spaces \cite{Tr78}.


\begin{lemma} \label{lem:vector-Lp-int}
Let $0<\theta<1$, $1\le p_0,p_1<\infty$ and $p\inv=(1-\theta)p_0\inv+\theta p_1\inv$.  Let $A_0,A_1$ be Banach spaces. Then 
\begin{equation}\label{eq:vector-Lp-int}
[L^{p_0}(A_0),L^{p_1}{(A_1)}]_\theta
=L^{p}([A_0,A_1]_\theta).   
\end{equation}
If $p_1=\infty$, 
then (\ref{eq:vector-Lp-int}) holds with 
$L^{p_1}(A_1)$ replaced by ${L_0^\infty}(A_1)$, the completion of simple $A_1$-valued functions with the esssup norm.
\end{lemma}

As in \cite{Tr78}, denote $\ell^q(A_j)$ the space of functions consisting of $a=\{a_j\}$, $a_j\in A_j$ ($A_j$ being Banach spaces)
equipped with the norm\\ 
$\displaystyle \Vert a\Vert_{\ell^q(A_j)}=\left( \sum_j\Vert a_j\Vert^q_{A_j} \right)^{1/q}$. 
\begin{lemma}\label{lem:vector-lq-int} Let  $0<\theta<1$, $1\le q_0,q_1<\infty$ and $q\inv=(1-\theta)q_0\inv+\theta q_1\inv$. 
Let $A_j$ be Banach spaces, $j\in\N$.
  Then \begin{equation}\label{e:ell-q-int}
[\ell^{q_0}(A_j),\ell^{q_1}(B_j)]_\theta
=\ell^{q}([A_j,B_j]_\theta).  \end{equation}
If 
$q_1=\iy$, then 
\begin{equation}\label{e:ell-q-infty-int}
[\ell^{q_0}(A_j),\ell^{\iy}(B_j)]_\theta
=\ell^{q}([A_j,B_j]_\theta)=[\ell^{q_0}(A_j),\ell_0^{\iy}(B_j)]_\theta  ,
\end{equation}
where $\ell_0^{\iy}(B_j):=\{ \{c_j\}\in \ell^{\iy}(B_j): \Vert c_j\Vert_{B_j}\to 0\;as\; j\to \iy\}$.
\end{lemma}

If $1\le q_0,q_1<\iy$, 
(\ref{e:ell-q-int}) also follows from 
Lemma \ref{lem:vector-Lp-int} as a special case where 
the underlying measure space can be taken as $(X,\mu)=\Z$. 
If $
q_1=\iy$, then the remark in \cite[Subsection 1.18.1]{Tr78}  shows that the second statement in 
Lemma \ref{lem:vector-lq-int} is also true.

 In the diagram above 
 in order to  show $S, R$ are continuous mappings,  we  need the following well-known lemma.


\begin{lemma}\label{lem:H-L-h} Let $h(x)$ be a monotonely nonincreasing, radial function 
in $L^1(\R^n)$. Let $h_j(x)=2^{jn/2}h(2^{j/2} x)$ be its scaling.  
Then for all $f$ in $L^1_{loc}(\R^n)$
\[
|\int h_j(x-y)f(y)dy|\le c_n\Vert h\Vert_1 Mf(x),
\]
where $Mf$ denotes the usual Hardy-Littlewood maximal function. 
\end{lemma}
Evidently the decay estimate in (\ref{e:ker-phi-dec}) and Lemma \ref{lem:H-L-h} imply the continuity of 
 $S$ and $R$,  in light of the $L^p(\ell^q)$-valued maximal inequality. 
The proof for $B^{s,q}_p(\cL)$ in part (b)  proceeds in a similar way. 

\subsection{Real interpolation} Peetre's $K$-functional \cite{Pe63} is defined as 
\begin{align*} K(t,a):=K(t,a;A_0,A_1)=\inf (\Vert a_0\Vert_{A_0}+t\Vert a_1\Vert_{A_1}) ,
\end{align*}
where the infimum is taken over all representations of $a=a_0+a_1$,  $a_i\in A_i$. 
Let $0<q\le\iy,0<\theta<1$.  For a given interpolation couple $(A_0,A_1)$, the real interpolation space $(A_0,A_1)_{\theta,q}$
is given by 
\begin{align*} 
(A_0,A_1)_{\theta,q}
=\{a\in A_0+A_1: \Vert a\Vert_{(A_0,A_1)_{\theta,q}}=\left( \int_0^\iy t^{-\theta q} K(t,a)^q \frac{dt}{t}\right)^{1/q}<\iy\}
\end{align*}
with usual modifications if $q=\iy$.


Proof of Theorem \ref{th:F-B-real-int} is similar to 
\cite[Subsection 2.4.2]{Tr83} and \cite[Theorem 6.4.5]{BL76}. 
Define $\ell^{s,q}(A)=\{a=\{a_j\}: \Vert a\Vert_{\ell^{s,q}(A)}=
\Vert \{2^{js}\Vert a_j\Vert_A\}\Vert_{\ell^q}<\iy\}$.
For Besov spaces it follows from  
\begin{align*}
&(\ell^{s_0,q_0}(A_0),\ell^{s_1,q_1}(A_1))_{\theta,q}=\ell^{s,q}((A_0,A_1)_{\theta,q}) ,\label{e:ell-sq-real} 
\end{align*}
 $s=(1-\theta)s_0+\theta s_1 $,
${q}\inv={(1-\theta)}{q_0\inv}+ {\theta}{q_1\inv}$  
 and the commutative diagram for $B_p^{s,q}(\cL)$. 
Consult  \cite{Tr83,Tr78} or
 \cite[Chapter 5,Theorem 6]{Pe76};  both of their proofs  rely on retraction method.
Also see \cite{BL76} for a different proof in the special case involving Sobolev spaces.
In the general case \cite{BL76}  suggests 
using a more concrete characterization of the $K$-functional for the Lorentz space $L^{pq}$.  

For the $F$-space 
  the proof follows from  the commutative diagram for $F_p^{s,q}(\cL)$ and
\begin{align*}
&(L^{p_0}(A_0,w_0),L^{p_1}(A_1,w_1))_{\theta,p}=L^{p}((A_0,A_1)_{\theta,p},w) ,
\end{align*}
where $ {p}\inv={(1-\theta)}{p_0\inv}+ {\theta}{ p_1\inv}$, $w=w_0^{1-\theta}w_1^\theta$,  
 $w_0,w_1$ being two weight functions 
 \cite[Chapter 5
 ]{Pe76}.

\subsection{Schr\"odinger operators with magnetic potential} 
From \cite
{He90a}, \cite{Z06a} 
or \cite{OZ08} 
we know that if the heat kernel of $\cL$ satisfies the upper Gaussian bound 
 \begin{equation}\label{e:g-b-etH}  |e^{-t\cL}(x,y)|\le c_n t^{-n/2} e^{-c |x-y|^2/t}
 \end{equation}
then 
the kernel decay in Assumption \ref{a:phi-dec-Om} holds. 
Let \begin{equation*}
H=  -\sum_{j=1}^n ( \partial_{x_j}+i a_j)^2+ V ,
\end{equation*}  where $a_j(x)\in L^2_{loc}(\R^n)$ is real-valued, 
$V=V_+-V_-$ with $V_+\in L^1_{loc}(\R^n)$, $V_-\in K_n$, the Kato class \cite{Si82}. 
 Proposition 5.1 in 
 \cite{DP05}  
 showed that (\ref{e:g-b-etH}) is valid  for $-\De+V$ 
 if $V_+\in K_n$ and 
$\Vert V_-\Vert_{K_n}< \ga_n:= \pi^{n/2}/\Ga(\frac{n}{2}-1)$, $n\ge 3$, 
whose proof evidently works for $V_+\in L^1_{loc}$. 
By the diamagnetic inequality  \cite[Theorem B.13.2]{Si82}, 
we see that \eqref{e:g-b-etH} also holds for $H$ provided 
$\Vert V_-\Vert_{K_n}< \ga_n$, $n\ge 3$. 

As another example,  a {\em uniformly 
elliptic operator} is given by 
\begin{equation*}
\cL=-\sum_{j,k=1}^n \pa_{x_j}(a_{jk}\pa_{x_k}) ,
\end{equation*}
where $a_{jk}(x)=a_{kj}(x)\in L^\iy(\R^n)$ are real-valued and satisfy the ellipticity condition $(a_{jk})\approx I_n$. 
Then \cite[Theorem 1]{ou06} tells that (\ref{e:g-b-etH}) is true provided that the infimum of its spectrum
$\inf\sigma (\cL)=0$.

\end{document}